# Provably Safe Model Updates

### v1

@@ I Introduction
Machine learning capabilities are evolving at unprecedented speed, driving deployment in high-stakes domains where safety and security are critical requirements [bommasani2021opportunities, szpruch2025insuring]. Models used in such critical domains (e.g., autonomous driving, medical applications, and financial services) must continually contend with input distribution shifts, newly discovered vulnerabilities, and evolving regulatory requirements. As a result, regular updates are a prerequisite to safe deployment. However, even benign updates can trigger severe unintended effects on model behavior—most notably catastrophic forgetting and loss of alignment—which are extensively documented in both foundation models [qi2023fine, yang2023shadow, zhan2023removing, lermen2023lora] and in continual learning [kirkpatrick2017overcoming, wang2024comprehensive]. In absence of a priori guarantees, model updates cannot be deployed into safety critical contexts without complete re-evaluation to satisfy regulators [us2024predetermined] or internal auditors [mokander2024auditing, wicker2025move]. Yet comprehensive re-validation is often impractical due to prohibitive computational cost, latency constraints, or inaccessible training data. This tension between the need for adaptation and the demand for assurance is a central obstacle to deploying learning systems safely in the real world.
Prior works usually seek to combat catastrophic forgetting or preserve model alignment during updates through approaches such as regularization [li2017learning, shen2020towards], replay buffers [rebuffi2017icarl, huang2024lisa], or parameter isolation [rusu2016progressive, kang2022forget]. While these methods display promising results, they do not come with formal guarantees of avoiding forgetting and therefore cannot avoid the need for revalidation. To address this, we investigate the problem of providing an update framework that has a priori guarantees that the updated model satisfies a given performance specification e.g., prior task performance or predictive stability. Beyond eliminating the revalidation bottleneck, we discuss how such a framework may offer certifiable safety evidence for regulators [us2024predetermined], reliable model version control, enforceable terms-of-use for downstream users [anderljung2023frontier, qi2024evaluating], and safe mechanisms for adapting to emerging vulnerabilities.
@@
The primary technical contribution of this paper is the development of effective and practical LID computation techniques as not all LIDs are practically useful. For example, the LID containing only the initial model parameter vector is trivially valid, but effectively prohibits model updates. Thus, we formulate the general LID problem as finding the maximal such domain (by volume) over all connected subsets of the parameter space. Because optimizing over arbitrary connected domains is intractable, we propose a tractable yet conservative relaxation: an optimization problem over parameterized abstract domains, which allows us to leverage advances in bounding neural networks [gowal2018effectiveness, tsay2021partition] and abstract interpretation [gehr2018ai2, sosnin2024certified]. Of course, the optimization problem over parameterized abstract domains must be constrained to only domains for which we can certify that every parameter respects the safety property. Thus, we instantiate a primal-dual formulation of the constrained maximal LID problem enabling the computation of approximately maximal LIDs via alternating gradient-descent-ascent.
@@
In Section V, we experimentally demonstrate the effectiveness of our approach across standard continual learning and foundation model fine-tuning tasks. We benchmark the performance of our algorithm against established regularization-based approaches and benchmark the guarantees of our algorithm against the continual learning approach InterContiNet (ICN) [wolczyk2022continual]. An embedded step of the ICN algorithm can be viewed as the computation of a LID, though the authors do not pose the problem as such. Thus, ICN does not consider finite sample error, use of replay buffers, LID multiplicity, or principled constraint encoding which leads to potentially trivial or unsound bounds and numerical instability. Across a wide variety of synthetic and real-world benchmarks including MNIST [lecun1998gradient], CIFAR-10 [CIFAR], hate speech classification, and multi-lingual adaptation, we find that our approach is able to match or best benchmark performance while providing state-of-the-art guarantees. In summary, this paper makes the following contributions:
@@ II Related Works
Ensuring alignment of foundation models is a growing area of research [huang2024harmful, zhaosurvey, huang2024survey]. Despite extensive training-time safeguards, it has been shown that fine-tuning rapidly erodes many safeguards [yang2023shadow], even when users are not malicious [qi2023fine]. Some heuristics have been developed to defend against this erosion [mukhoti2023fine, wei2024assessing, rosati2024representation], but have been found unsuccessful in the face of more sophisticated attackers [qi2024evaluating, rando2025adversarial]. These studies indicate the start of an arms race between attackers and defenders often observed in adversarial settings, and they point to the importance of formal verification in safety-critical contexts.
@@
Ensuring that training updates do not degrade prior model capabilities (as seen in Figure 1), an issue termed catastrophic forgetting [kirkpatrick2017overcoming, li2017learning, van2022three], has been widely studied within the continual learning community [wang2024comprehensive]. Avoiding forgetting is often studied under the assumption that data from prior tasks are unavailable [van2022three]. In this case, heuristic forgetting defenses can largely be split into approaches that use parameter isolation (i.e., only updating a subset of parameters [rusu2016progressive, yoon2017lifelong, kang2022forget]) or regularization [li2017learning, kirkpatrick2017overcoming, shen2020towards] (i.e., soft constraints limiting changes to important parameters). With access to prior task data for revalidation, replay methods have been developed [rebuffi2017icarl, huang2024lisa]. Each of these methods has demonstrated some considerable success in reducing forgetting; however, they do not come with concrete, formal guarantees.
@@
The field of formal methods for machine learning algorithms has produced proofs of a wide range of notions including adversarial robustness [gowal2018effectiveness, sosnin2024certified], explainability [wicker2022robust], privacy [wicker2024certification], and uncertainty [bitterwolf2020certifiably]. Most related to this paper is the growing literature on formal security guarantees for machine learning models [wang2022improved, sosnin2024certified] which, however, focus on specific dataset threat models. The most closely related work is InterContiNet (ICN) [wolczyk2022continual], which introduces a continual learning algorithm that uses heuristic optimization to identify safe parameter intervals. The algorithm begins with a large, potentially unsafe parameter interval and iteratively contracts it via SGD to meet an accuracy threshold. In contrast to our approach, ICN seeks only to find some valid LID rather than a maximal LID, reducing the model’s flexibility to learn new tasks while preserving safety. Moreover, its shrinkage-based optimization procedure is susceptible to interval collapse and numerical instability, which hinders scalability to larger models. Finally, the method does not incorporate multiplicity, biasing, replay buffers, or finite-sample safety guarantees – key contributions of this work.
@@ III Certification for Provably Safe Updates
Throughout the paper we will consider tasks to be supervised learning problem with a fixed sample size $N$, i.e., $\mathcal{D}:=\{(x^{(i)},y^{(i)}\}_{i=1}^{N}$, where we describe $x\in\mathbb{R}^{n}$ as our feature vectors and $y\in\mathcal{Y}$ to be our labels; we will assume labels inhabit a $c$-dimensional space (i.e., $c$ classes or $c$-dimensional regression). We denote machine learning models as parameterized functions $f^{\theta}:\mathbb{R}^{n}\to\mathcal{Y}$ with parameters $\theta\in\mathbb{R}^{p}$. A proposed update to the parameters comprises a vector $u\in\mathbb{R}^{p}$, with the result of applying the update being: $\theta^{\prime}=\theta+u$. We denote a function that proposes model updates (SGD, SFT, DPO, etc.) as an update mechanism $\mathcal{M}$. Finally, we model behavioral properties of interest such as alignment or performance criteria using a function $\phi:\mathbb{R}^{p\times n\times c}\to\mathbb{R}$. Without loss of generality, we use the convention that lower scores of $\phi$ are better and indicate better performance/alignment, and in our experiments we take $\phi$ to be one minus the accuracy on a particular task, but this is left purposefully generic with further discussion of concrete choices in Appendix -C. We now provide the general problem statement for a provably safe update mechanism:
@@
($\delta$-Safe Update Mechanism) Given a machine learning model $f^{\theta}$ trained on a task with data distribution $P(X,Y)$ and a property $\phi$ we say that a mechanism $\mathcal{M}$ taking in an arbitrary dataset $\mathcal{D}^{\prime}$ and parameter $\theta$ is a provably $\delta$-safe update mechanism if:
@@ III-A Local Invariant Domains (LIDs)
(Locally Invariant Domain) Given a machine learning model $f^{\theta}$ trained on a task with data distribution $P(X,Y)$, and a property $\phi$, we say that a connected domain $T\subseteq\mathbb{R}^{p}$ containing $\theta$ is a provably $\delta$-invariant domain if:
@@ III-B2 Multi-Step Updates (Continual Learning)
In continual learning, the model is updated multiple times [van2022three], and each update must ensure that all previous tasks respect a given constraint. Definition 1 can be generalized to $k$-many, ordered updates by defining the $k^{th}$ update safety as: $\forall\mathcal{D}^{\prime}_{k},\forall i\in[k-1],$
@@ III-B3 Use of Replay Buffer
$\forall\theta^{\prime}\in T,i\in[k-1],\,\mathbb{E}_{P(X_{i},Y_{i})},\Big[\phi\big(\theta^{\prime},x,y\big)\Big]\leq\delta_{i}$
@@ III-B4 LID Biasing
Despite formalizing the maximal locally invariant domain problem in Definition 3, the maximization is posed with respect to current task. Even if one computes the exact maximal LID, the domain may not span directions that are important for improving performance on subsequent tasks. Identifying such important directions has been the focus of many prior continual learning approaches. In particular, we draw from weight pruning literature, which uses weight importance measures to determine parameters that significantly contribute to the performance of a given model [Mallya2018PackNet, golkar2019CLNP]. Below, we discuss how these importance methods, broken down into weight pruning or regularization, can be used to bias LID computation.
Weight pruning approaches define a discrete set of weights that are determined to be important. Formally, let $\mathcal{I}(\cdot)$ be a function associated with a weight pruning method that maps parameter indices to either 0 (if non-important) or 1 (if important) as determined by the pruning method. We can then restrict our prior optimization formulation to only be over only the unimportant parameter subspace i.e.., leaving all important weights fixed. Where we denote members of this subspace with $\mathbb{R}^{p}_{\mathcal{I}}\subseteq\mathbb{R}^{p}$ we can reformulate the optimization problem as:
$\max_{T\in\mathbb{R}^{p}_{\mathcal{I}}}|T|_{S}\quad s.t.\quad\max_{\theta^{\prime}\in T}\left[\frac{1}{N}\sum_{i=1}^{N}\phi(\theta^{\prime},x_{i},y_{i})\right]\leq\delta$
Intuitively, this modification only allows for updates to those weights that are unimportant for the current task, making it less likely that one violates the stated constraint (after $s.t.$ above). A critical parameter in weight pruning methods is the proportion, $\alpha\in(0,1)$ of weights to be pruned – we explore this hyper-parameter in detail in Section V.
@@
$\max_{T\in\mathbb{R}^{p}}|T|_{S}+\mathcal{R}(T)\quad s.t.\quad\max_{\theta^{\prime}\in T}\left[\frac{1}{N}\sum_{i=1}^{N}\phi(\theta^{\prime},x_{i},y_{i})\right]\leq\delta$
@@ III-B5 Finite-Sample Guarantees for $\delta$-Safe Updates
Definition 1 involves a constraint with an intractable integral over the distribution $P(X,Y)$, which is typically taken to be unknown. In practice, we approximate the expectation using a finite evaluation set that is held out during training (or held out within a data buffer): $\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim P(X,Y)$, with the empirical estimate denoted as,
@@ IV-A Propagation of Abstract Domains
The inner maximization in (4) corresponds to a parameter-space verification problem, which seeks to upper bound the empirical loss over a set of model parameter values. This formulation is analogous to recent work on verifying robustness to parameter-space perturbations [wicker2022robust, sosnin2024certified], but differs from standard adversarial robustness settings that consider perturbations over the input space. Specifically, we are concerned with the worst-case model outputs over all parameter vectors rather than input vectors. For ease of exposition, we discuss the case where $T$ is the interval domain, denoted $T(\alpha)$ with $\alpha:=[\alpha^{L},\alpha^{U}]$, which encodes all parameter vectors $\theta\in\mathbb{R}^{p}$ satisfying
@@
This representation defines a box in parameter space and serves as an abstract domain, formally known as the orthotope domain, over which we can compute bounds on the output of a neural network using Interval Bound Propagation (IBP) [gowal2018effectiveness].
Given a neural network $f^{\theta}$ and an interval domain $T(\alpha)$, IBP computes lower and upper bounds on the network output at an input point $x$ for any $\theta\in T(\alpha)$. This propagation relies on all used activation functions being monotonic (e.g., ReLU, tanh) [gowal2018effectiveness]. This assumption should not be substantially restrictive in practice, as most commonly used activation functions are monotonic. Moreover, there are some ways to relax this requirement. Subsequently, these output bounds can then be propagated through the specification function $\phi$ to obtain a sound upper bound on the empirical specification estimate. In particular, we use the following IBP-based relaxation:
@@ IV-B Primal-Dual Optimization for Computing Maximal Locally Invariant Domains
Interval domains offer not only a natural parameterization for the constrained problem in (4), but also a differentiable and sound over-approximation of the constraint. This yields a single-level constrained optimization problem that constitutes a sound under-approximation of the original formulation. That is, any solution identified via this relaxation defines a LID that satisfies the safety specification, although it may be strictly smaller than the solution of the original bi-level problem (4).
@@
This gives rise to the saddle-point optimization $\max_{\alpha}\min_{\lambda\geq 0}L(\alpha,\lambda)$ which can be locally solved using alternating gradient descent-ascent [zhang2022near], which performs the updates:
@@
where $\eta_{d}$ and $\eta_{p}$ are step sizes for the dual and primal updates, respectively, and $\cdot$$_{+}$ denotes projection onto the non-negative orthant to ensure $\lambda\geq 0$ [^2] We employ the cooper library for Lagrangian-based optimization [gallego2025cooper].. We additionally apply an extra projection step to $\alpha$ that enforces that the nominal parameters $\theta$ remain inside $T(\alpha)$. The above primal-dual optimization procedure is designed to maintain feasibility with respect to the safety specification constraint throughout training. However, due to the inherent stochasticity of gradient-based optimization, the final iterate may not strictly satisfy the original constraint. To formally verify that the returned LID, denoted $T(\alpha)$, satisfies the desired specification, we compute a post-optimization certificate by computing $\max_{\theta^{\prime}\in T(\alpha)}\frac{1}{n}\sum_{i=1}^{n}\phi(\theta^{\prime},x_{i},y_{i})$.

### v2

@@ I Introduction
Machine learning capabilities are evolving at unprecedented speed, driving deployment in high-stakes domains where safety and security are critical requirements [6, 45]. Models used in such critical domains (e.g., autonomous driving, medical applications, and financial services) must continually contend with input distribution shifts, newly discovered vulnerabilities, and evolving regulatory requirements. As a result, regular updates are a prerequisite to safe deployment. However, even benign updates can trigger severe unintended effects on model behavior—most notably catastrophic forgetting and loss of alignment—which are extensively documented in both foundation models [35, 59, 63, 28] and in continual learning [24, 49]. In absence of a priori guarantees, model updates cannot be deployed into safety critical contexts without complete re-evaluation to satisfy regulators [47] or internal auditors [31, 57]. Yet comprehensive re-validation is often impractical due to prohibitive computational cost, latency constraints, or inaccessible training data. This tension between the need for adaptation and the demand for assurance is a central obstacle to deploying learning systems safely in the real world.
Prior works usually seek to combat catastrophic forgetting or preserve model alignment during updates through approaches such as regularization [29, 43], replay buffers [37, 19], or parameter isolation [41, 21]. While these methods display promising results, they do not come with formal guarantees of avoiding forgetting and therefore cannot avoid the need for revalidation. To address this, we investigate the problem of providing an update framework that has a priori guarantees that the updated model satisfies a given performance specification e.g., prior task performance or predictive stability. Beyond eliminating the revalidation bottleneck, we discuss how such a framework may offer certifiable safety evidence for regulators [47], reliable model version control, enforceable terms-of-use for downstream users [3, 34], and safe mechanisms for adapting to emerging vulnerabilities.
@@
The primary technical contribution of this paper is the development of effective and practical LID computation techniques as not all LIDs are practically useful. For example, the LID containing only the initial model parameter vector is trivially valid, but effectively prohibits model updates. Thus, we formulate the general LID problem as finding the maximal such domain (by volume) over all connected subsets of the parameter space. Because optimizing over arbitrary connected domains is intractable, we propose a tractable yet conservative relaxation: an optimization problem over parameterized abstract domains, which allows us to leverage advances in bounding neural networks [17, 46] and abstract interpretation [15, 44]. Of course, the optimization problem over parameterized abstract domains must be constrained to only domains for which we can certify that every parameter respects the safety property. Thus, we instantiate a primal-dual formulation of the constrained maximal LID problem enabling the computation of approximately maximal LIDs via alternating gradient-descent-ascent.
@@
In Section V, we experimentally demonstrate the effectiveness of our approach across standard continual learning and foundation model fine-tuning tasks. We benchmark the performance of our algorithm against established regularization-based approaches and benchmark the guarantees of our algorithm against the continual learning approach InterContiNet (ICN) [58]. An embedded step of the ICN algorithm can be viewed as the computation of a LID, though the authors do not pose the problem as such. Thus, ICN does not consider finite sample error, use of replay buffers, LID multiplicity, or principled constraint encoding which leads to potentially trivial or unsound bounds and numerical instability. Across a wide variety of synthetic and real-world benchmarks including MNIST [27], CIFAR-10 [25], hate speech classification, and multi-lingual adaptation, we find that our approach is able to match or best benchmark performance while providing state-of-the-art guarantees. In summary, this paper makes the following contributions:
@@ II Related Works
Ensuring alignment of foundation models is a growing area of research [18, 65, 20]. Despite extensive training-time safeguards, it has been shown that fine-tuning rapidly erodes many safeguards [59], even when users are not malicious [35]. Some heuristics have been developed to defend against this erosion [32, 51, 38], but have been found unsuccessful in the face of more sophisticated attackers [34, 36]. These studies indicate the start of an arms race between attackers and defenders often observed in adversarial settings, and they point to the importance of formal verification in safety-critical contexts.
@@
Ensuring that training updates do not degrade prior model capabilities (as seen in Figure 1), an issue termed catastrophic forgetting [24, 29, 48], has been widely studied within the continual learning community [49]. Avoiding forgetting is often studied under the assumption that data from prior tasks are unavailable [48]. In this case, heuristic forgetting defenses can largely be split into approaches that use parameter isolation (i.e., only updating a subset of parameters [41, 61, 21]) or regularization [29, 24, 43] (i.e., soft constraints limiting changes to important parameters). With access to prior task data for revalidation, replay methods have been developed [37, 19]. Each of these methods has demonstrated some considerable success in reducing forgetting; however, they do not come with concrete, formal guarantees.
@@
The field of formal methods for machine learning algorithms has produced proofs of a wide range of notions including adversarial robustness [17, 44], explainability [52], privacy [56], and uncertainty [5]. Most related to this paper is the growing literature on formal security guarantees for machine learning models [50, 44] which, however, focus on specific dataset threat models. The most closely related work is InterContiNet (ICN) [58], which introduces a continual learning algorithm that uses heuristic optimization to identify safe parameter intervals. The algorithm begins with a large, potentially unsafe parameter interval and iteratively contracts it via SGD to meet an accuracy threshold. In contrast to our approach, ICN seeks only to find some valid LID rather than a maximal LID, reducing the model’s flexibility to learn new tasks while preserving safety. Moreover, its shrinkage-based optimization procedure is susceptible to interval collapse and numerical instability, which hinders scalability to larger models. Finally, the method does not incorporate multiplicity, biasing, replay buffers, or finite-sample safety guarantees—key contributions of this work.
@@ III Certification for Provably Safe Updates
Throughout the paper we will consider tasks to be supervised learning problem with a fixed sample size $N$, i.e., $\mathcal{D}:=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}$, where we describe $x\in\mathbb{R}^{n}$ as our feature vectors and $y\in\mathcal{Y}$ to be our labels; we will assume labels inhabit a $c$-dimensional space (i.e., $c$ classes or $c$-dimensional regression). We denote machine learning models as parameterized functions $f^{\theta}:\mathbb{R}^{n}\to\mathcal{Y}$ with parameters $\theta\in\mathbb{R}^{p}$. A proposed update to the parameters comprises a vector $u\in\mathbb{R}^{p}$, with the result of applying the update being: $\theta^{\prime}=\theta+u$. We denote a function that proposes model updates (SGD, SFT, DPO, etc.) as an update mechanism $\mathcal{M}$. Finally, we model behavioral properties of interest such as alignment or performance criteria using a function $\phi:\mathbb{R}^{p\times n\times c}\to\mathbb{R}$. Without loss of generality, we use the convention that lower scores of $\phi$ are better and indicate better performance/alignment, and in our experiments we take $\phi$ to be one minus the accuracy on a particular task, but this is left purposefully generic with further discussion of concrete choices in Appendix -C, alongside a collection of notation used throughout this section in Appendix -H. We now provide the general problem statement for a provably safe update mechanism:
@@
($\delta$-Safe Update Mechanism) Given a machine learning model $f^{\theta}$ trained on a task with data distribution $P(X,Y)$ and a property $\phi$ we say that a mechanism $\mathcal{M}$ taking in an arbitrary dataset $\mathcal{D}^{\prime}$ and parameter $\theta$ is a provably $\delta$-safe update mechanism if
@@ III-A Local Invariant Domains (LIDs)
(Locally Invariant Domain) Given a machine learning model $f^{\theta}$ trained on a task with data distribution $P(X,Y)$, and a property $\phi$, we say that a connected domain $T\subseteq\mathbb{R}^{p}$ containing $\theta$ is a provably $\delta$-invariant domain if
@@ III-B2 Multi-Step Updates (Continual Learning)
In continual learning, the model is updated multiple times [48], and each update must ensure that all previous tasks respect a given constraint. Definition 1 can be generalized to $k$-many, ordered updates by defining the $k^{th}$ update safety as: $\forall\mathcal{D}^{\prime}_{k},\forall i\in[k-1],$
@@ III-B3 Use of Replay Buffer
$\forall\theta^{\prime}\in T,i\in[k-1],\,\mathbb{E}_{P(X_{i},Y_{i})},\Big[\phi\big(\theta^{\prime},x,y\big)\Big]\leq\delta_{i},$
@@ III-B4 LID Biasing
Despite formalizing the maximal locally invariant domain problem in Definition 3, the maximization is posed with respect to current task. Even if one computes the exact maximal LID, the domain may not span directions that are important for improving performance on subsequent tasks. Identifying such important directions has been the focus of many prior continual learning approaches. In particular, we draw from weight pruning literature, which uses weight importance measures to determine parameters that significantly contribute to the performance of a given model [30, 16]. Below, we discuss how these importance methods, broken down into weight pruning or regularization, can be used to bias LID computation.
Weight pruning approaches define a discrete set of weights that are determined to be important. Formally, let $\mathcal{I}(\cdot)$ be a function associated with a weight pruning method that maps parameter indices to either 0 (if non-important) or 1 (if important) as determined by the pruning method. We can then restrict our prior optimization formulation to only be over only the unimportant parameter subspace i.e.., leaving all important weights fixed. Where we denote members of this subspace with $\mathbb{R}^{p}_{\mathcal{I}}\subseteq\mathbb{R}^{p}$ we can reformulate the optimization problem as
$\max_{T\in\mathbb{R}^{p}_{\mathcal{I}}}|T|_{S}\quad s.t.\quad\max_{\theta^{\prime}\in T}\left[\frac{1}{N}\sum_{i=1}^{N}\phi(\theta^{\prime},x_{i},y_{i})\right]\leq\delta.$
Intuitively, this modification only allows for updates to those weights that are unimportant for the current task, making it less likely that one violates the stated constraint (after $s.t.$ above). A critical parameter in weight pruning methods is the proportion, $\alpha\in(0,1)$ of weights to be pruned—we explore this hyper-parameter in detail in Section V.
@@
$\max_{T\in\mathbb{R}^{p}}|T|_{S}+\mathcal{R}(T)\quad s.t.\quad\max_{\theta^{\prime}\in T}\left[\frac{1}{N}\sum_{i=1}^{N}\phi(\theta^{\prime},x_{i},y_{i})\right]\leq\delta.$
@@ III-B5 Finite-Sample Guarantees for $\delta$-Safe Updates
Definition 1 involves a constraint with an intractable integral over the distribution $P(X,Y)$, which is typically taken to be unknown. In practice, we approximate the expectation using a finite evaluation set that is held out during training (or held out within a data buffer): $\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim P(X,Y)$, with the empirical estimate denoted as
@@ IV-A Propagation of Abstract Domains
The inner maximization in (4) corresponds to a parameter-space verification problem, which seeks to upper bound the empirical loss over a set of model parameter values. This formulation is analogous to recent work on verifying robustness to parameter-space perturbations [52, 44], but differs from standard adversarial robustness settings that consider perturbations over the input space. Specifically, we are concerned with the worst-case model outputs over all parameter vectors rather than input vectors. For ease of exposition, we discuss the case where $T$ is the interval domain, denoted $T(\alpha)$ with $\alpha:=[\alpha^{L},\alpha^{U}]$, which encodes all parameter vectors $\theta\in\mathbb{R}^{p}$ satisfying
@@
This representation defines a box in parameter space and serves as an abstract domain, formally known as the orthotope domain, over which we can compute bounds on the output of a neural network using Interval Bound Propagation (IBP) [17].
Given a neural network $f^{\theta}$ and an interval domain $T(\alpha)$, IBP computes lower and upper bounds on the network output at an input point $x$ for any $\theta\in T(\alpha)$. This propagation relies on all used activation functions being monotonic (e.g., ReLU, tanh) [17]. This assumption should not be substantially restrictive in practice, as most commonly used activation functions are monotonic. Moreover, there are some ways to relax this requirement. Subsequently, these output bounds can then be propagated through the specification function $\phi$ to obtain a sound upper bound on the empirical specification estimate. In particular, we use the following IBP-based relaxation:
@@ IV-B Primal-Dual Optimization for Computing Maximal Locally Invariant Domains
Interval domains offer not only a natural parameterization for the constrained problem in Equation  (4), but also a differentiable and sound over-approximation of the constraint. This yields a single-level constrained optimization problem that constitutes a sound under-approximation of the original formulation. That is, any solution identified via this relaxation defines a LID that satisfies the safety specification, although it may be strictly smaller than the solution of the original bi-level problem in Equation  (4).
@@
This gives rise to the saddle-point optimization $\max_{\alpha}\min_{\lambda\geq 0}L(\alpha,\lambda)$ which can be locally solved using alternating gradient descent-ascent [64], which performs the updates:
@@
where $\eta_{d}$ and $\eta_{p}$ are step sizes for the dual and primal updates, respectively, and $\cdot$$_{+}$ denotes projection onto the non-negative orthant to ensure $\lambda\geq 0$ [^2] We employ the cooper library for Lagrangian-based optimization [14].. We additionally apply an extra projection step to $\alpha$ that enforces that the nominal parameters $\theta$ remain inside $T(\alpha)$. The above primal-dual optimization procedure is designed to maintain feasibility with respect to the safety specification constraint throughout training. However, due to the inherent stochasticity of gradient-based optimization, the final iterate may not strictly satisfy the original constraint. To formally verify that the returned LID, denoted $T(\alpha)$, satisfies the desired specification, we compute a post-optimization certificate by computing $\max_{\theta^{\prime}\in T(\alpha)}\frac{1}{n}\sum_{i=1}^{n}\phi(\theta^{\prime},x_{i},y_{i})$.